A vapour power plant working on Rankine cycle operates between boiler pressure of 500KN/M2 and condenser pressure of 0.06 bar calculate (1) The Thermal Efficiency (2) Work ratio (3) Specific steam consumption for the cycle.


1
Expert's answer
2021-06-22T05:10:00-0400

Form the tables "h_1= 2747.5; h_2= 151.5+0.88*2416=2277.59; h_3= 151.5; h_4=640"


(1) The Thermal Efficiency

"η_T= \frac{h_1-h_2}{h_1-h_4}*100= \frac{2747.5-2277.59}{2747.5-640}*100= 22 \%"


(2) Work ratio

Work ratio "=1-\frac{W_P}{W_T}=1-\frac{h_4-h_3}{h_1-h_4}=1-\frac{640-151.5}{2747.5-640}=0.7681"


(3) Specific steam consumption for the cycle.

"SSC = \frac{3600}{W_{net}} = \frac{3600}{W_T-W_P}= \frac{3600}{2107-488} = \frac{3600}{1618.5}=2.22 kg/kWh"